# A Prompt-Guided Spatio-Temporal Transformer Model for National-Wide Nuclear Radiation Forecasting

### v1

@@ 5. Experiments
In this section, we introduce the experiments setup, overall performance, ablation study and parameter sensitivity analysis. Due to the page limit, we report the details of the web-based system design to Appendix C.
@@ 5.1.1. Evaluation Metrics
We employed three widely used metrics, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE), for model evaluation. Lower values in MAE, RMSE, and MAPE indicate higher forecasting accuracy. Moreover, we follow (Zheng et al., 2015; Yi et al., 2018; Liang et al., 2023) to discuss the errors on predicting sudden changes. We identified the top 10% of samples in the test set with the greatest variation between observed and predicted values as sudden change samples for prediction. We provide detailed information on evaluation metrics in Appendix B.2.
@@ 5.1.2. Baseline
We compare the proposed NRFormer with 11 baseline methods, including Historical Average (HA) (Zhang et al., 2017), XGBoost (Chen and Guestrin, 2016), Linear Regression (LR), DCRNN (Li et al., 2018), Graph WaveNet (Wu et al., 2019), STID (Shao et al., 2022b), DLinear (Zeng et al., 2023), StemGNN (Cao et al., 2020), LightCTS (Lai et al., 2023), PatchTST (Nie et al., 2023) and Koopa (Liu et al., 2024b). We provide detailed baselines in Appendix B.1.
@@ 5.1.3. Implementation Details
The dataset was partitioned into three distinct subsets: 60% for training, 20% for validation, and 20% for testing. Our experimens are conducted on two different server configurations. The first configuration consisted of three Linux Centos servers, each equipped with four RTX 3090 GPUs. In contrast, the second configuration utilized a single Linux Ubuntu server with two A800 GPUs. In terms of implementation, our model is implemented using PyTorch, with the Adam optimizer chosen for optimization. We set the learning rate and the weight decay to 0.001 and 0.0001, respectively. The batch size is fixed to 32. The number of temporal and spatial attention layers are fixed to 3, and the head of the multi-head attention module is set to 4.
@@ 5.2. Overall Performance
Table 2 presents the overall performance of NRFormer and all the baseline models on two real-world datasets with respect to MAE, RMSE, and MAPE. Overall, NRFormer consistently outperforms other baseline models on both datasets, demonstrating the effectiveness of the proposed model. Specifically, our model achieves improvements of 8.02%, 10.95%, 13.95%, and 13.1% beyond the best baseline (LightCTS) in terms of the MAE metric on 6, 9, 12, and 24 steps on the Japan-4H dataset, respectively. Similarly, the improvement of the MAE on the Japan-1D dataset is 7.57%, 7.77%, 8.5%, and 6.51%, respectively. Moreover, we can make the following observations: (1) All traditional methods (i.e., HA, LR, and XGBoost) perform worse than deep learning models, as they fail to capture spatio-temporal patterns. (2) STGNN-based approaches (i.e., DCRNN, Graph WaveNet) outperform HA, LR, and XGBoost by a large margin, indicating the superiority of modeling intricate spatio-temporal dependencies. (3) STID performs worse than DCRNN and Graph WaveNet on Japan-1D dataset. The possible reason is that Japan-1D has fewer training samples than Japan-4H, which is insufficient to learn discriminative spatial and temporal identity embeddings. (4) LightCTS leverages a tailored designed automatic spatio-temporal feature extraction module and exhibits strong learning capabilities, achieving the best performance compared with other baseline models.
@@
5.3. Ablation Study
We also examine the utility of the contextual features we provide to the prompting module. The results on MAE and RMSE are presented in Figure 6. In the absence of any contextual features, we separately add one feature at each time to assess its contribution. Here noPrompt indicates the performance of NRFormer without utilizing any additional contextual features. We can observe that incorporating each feature consistently improves the model performance beyond the noPrompt. In summary, the contextual features have a significant impact on nuclear radiation, and taking them into account is beneficial for forecasting.
Figure 6. Ablation study of each contextual features on Japan-4H and Japan-1D datasets. noPrompt means that we remove all the contextual features from our framework. Others (e.g., Wind Speed) indicate using only one contextual feature.
@@ 5.3. Ablation Study
5.4. Parameter Sensitivity Analysis
6. Related Works
Transformer. The advent of the Transformer architecture (Vaswani et al., 2017) has been making a transformative impact across various fields, particularly in time series forecasting (Zhou et al., 2021; Cao et al., 2023; Chowdhury et al., 2022; Gorbett et al., 2023) and graph mining (Zhang et al., 2021; Zhu et al., 2023; Fan et al., 2021; Li et al., 2022; Wu et al., 2022). Transformer depends on self-attention mechanism to capture data correlations and is capable of modeling complicated and long-range dependencies. Due to the powerful capability, recently considerable endeavors have been made to adapt Transformer to time series and spatio-temporal data (Zha et al., 2021; Feng and Tassiulas, 2022; Liang et al., 2023; Zhang and Yan, 2022; Liu et al., 2023). For instance, by introducing a ProbSparse self-attention mechanism, Informer (Zhou et al., 2021) enhances the Transformer framework for time series forecasting with high efficiency and scalability. As another example, Autoformer(Wu et al., 2021) further pushes the boundaries of time series forecasting by integrating decomposition and auto-correlation mechanism into Transformer. In this paper, we adapt the Transformer model with tailored designs to address the unique challenges in nuclear radiation forecasting.
Figure 7. Parameter sensitivity of NRFormer.
@@ 7. Conclusion
In this paper, we introduce NRFormer, a spatio-temporal graph Transformer for national-wide nuclear radiation forecasting. Specifically, we first devise a non-stationary temporal attention module to capture the erratic and non-stationary characteristics in radiation time series. Subsequently, we propose an imbalance-aware spatial attention module to address the imbalanced spatial distribution issue by adaptively re-weighting the influence of stations from both global and local perspectives. Furthermore, a radiation propagation prompting module is developed to guide the predictive modeling process. Extensive experiments on real-world nuclear radiation datasets demonstrate the superiority of NRFormer against seven baselines. In the future, we plan to deploy NRFormer to more countries so as to empower various decision-making tasks such as emergency response planning, thereby protecting environmental safety and public health.
Appendix A Relevance To Web
The increasing prevalence of the Web of Things (WoT) has facilitated the continuous collection of spatiotemporal data from sensors deployed across various geographic locations. Spatiotemporal modeling (Yang et al., 2024; Long et al., 2024; Zhang et al., 2023; Zheng et al., 2023) has been widely utilized in web applications, such as social networks (Miyake et al., 2024; Zhou et al., 2023) and Web of Things applications (Guo et al., 2024; He et al., 2022; He and Shin, 2022). Radiation forecasting is a typical spatiotemporal task within the domains of web mining and knowledge discovery, playing a crucial role in aggregating and analyzing real-time radiation data. This allows seamless access to critical information for timely decision-making. We developed a web-based system to provide comprehensive, nationwide predictions of nuclear radiation levels in Japan. This system helps people understand future radiation levels in their surroundings, enabling accurate risk assessments and early disaster warnings, ultimately safeguarding public health and environmental safety.
Appendix B Experimental Setup
B.1. Baselines
@@
  • DCRNN (Li et al., 2018): a spatio-temporal forecasting model that captures spatial dependencies through bidirectional random walks on the graph, while addressing temporal dependencies using an encoder-decoder architecture with scheduled sampling.
@@
  • PatchTST (Nie et al., 2023): PatchTST is an efficient Transformer-based model for multivariate time series forecasting, leveraging subseries patching and channel independence to reduce complexity and improve long-term accuracy, also excelling in self-supervised tasks and setting new accuracy benchmarks.
  • Koopa (Liu et al., 2024b): Koopa introduces a novel approach for forecasting non-stationary time series by leveraging modern Koopman theory to disentangle and advance time-variant and time-invariant dynamics.
B.2. Evaluation Metrics
We utilize four commonly adopted metrics to evaluate the performance of the model, namely Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are defined as follows:
(21) $MAE=\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}|\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij}|,$
(22) $RMSE=\sqrt{\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}(\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij})^{2}},$
@@ B.2. Evaluation Metrics
Appendix C Web-based System
We developed a web-based system to provide comprehensive, nationwide prediction of nuclear radiation levels in Japan. Specifically, we provide a heatmap of daily radiation forecasts across all of Japan. As can be seen, the heatmap shows the detailed spatial variation of radiation forecasts, enabling a comprehensive understanding of how nuclear radiation spreads across different regions of the country over time.
This web-based system also allows users to track the future changes in radiation levels over time. In this system, each blue point on the map represents a specific monitoring station. By clicking a station on the map, a pop-up chart appears, displaying a curve that forecasts radiation levels for the upcoming 24 days. Users can freely choose any monitoring stations at a particular time. This visualization provides valuable information on future radiation trends, which can aid in various decision-making processes about environmental and public risk management.

### v2

@@ 5. Experiments
In this section, we introduce the experiment setup, overall performance, ablation study, and deployed system evaluation. Due to the page limit, we report the details of parameter sensitivity analysis in Appendix A.3.
@@ 5.1. Experiments Setup
5.1.1. Datasets
The experimental datasets for this predictive model include nuclear radiation data and meteorological data collected from the entire country of Japan. Table 1 summarizes the statistics of the radiation and meteorological datasets.
@@ 5.1.2. Evaluation Metrics
We employed three widely used metrics, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE), for model evaluation. Lower values in MAE, RMSE, and MAPE indicate higher forecasting accuracy. Moreover, we follow (Zheng et al., 2015; Yi et al., 2018; Liang et al., 2023) to discuss the errors on predicting sudden changes. We identified the top 10% of samples in the test set with the greatest variation between observed and predicted values as sudden change samples for prediction. We provide detailed information on evaluation metrics in Appendix A.2.
@@ 5.1.3. Baseline
We compare the proposed NRFormer with 11 baseline methods, including Historical Average (HA) (Zhang et al., 2017), XGBoost (Chen and Guestrin, 2016), Linear Regression (LR), DCRNN (Li et al., 2018), Graph WaveNet (Wu et al., 2019), STID (Shao et al., 2022b), DLinear (Zeng et al., 2023), StemGNN (Cao et al., 2020), LightCTS (Lai et al., 2023), PatchTST (Nie et al., 2023) and Koopa (Liu et al., 2024). We provide detailed baselines in Appendix A.1.
@@ 5.1.4. Implementation Details
The dataset was partitioned into three distinct subsets: 60% for training, 20% for validation, and 20% for testing. Our experimens are conducted on two different server configurations: Linux Centos with four RTX 3090 GPUs and Linux Ubuntu with two A800 GPUs. In terms of implementation, our model is implemented using PyTorch, with the Adam optimizer chosen for optimization. We set the learning rate and the weight decay to 0.001 and 0.0001, respectively. The batch size is fixed to 32. The number of temporal and spatial attention layers are fixed to 3, and the head of the multi-head attention module is set to 4.
@@ 5.2. Overall Performance
Table 2 presents the overall performance of NRFormer and all the baseline models on two real-world datasets with respect to MAE, RMSE, and MAPE. Overall, NRFormer consistently outperforms other baseline models on both datasets, demonstrating the effectiveness of NRFormer. Specifically, our model achieves improvements of 8.02%, 10.95%, 13.95%, and 13.1% beyond the LightCTS in terms of the MAE metric on 6, 9, 12, and 24 steps on the Japan-4H dataset, respectively. Similarly, the improvement of the MAE on the Japan-1D dataset is 7.57%, 7.77%, 8.5%, and 6.51%, respectively. Moreover, we can make the following observations: (1) All traditional methods (i.e., HA, LR, and XGBoost) perform worse than deep learning models, as they fail to capture spatio-temporal patterns. (2) STGNN-based approaches (i.e., DCRNN, Graph WaveNet) outperform HA, LR, and XGBoost by a large margin, indicating the superiority of modeling intricate spatio-temporal dependencies. (3) STID performs worse than DCRNN and Graph WaveNet on Japan-1D dataset. The possible reason is that Japan-1D has fewer training samples than Japan-4H, which is insufficient to learn discriminative spatial and temporal identity embeddings. (4) LightCTS leverages a tailored designed automatic spatio-temporal feature extraction module and exhibits strong learning capabilities, achieving the best performance compared with other baseline models.
@@
5.3. Ablation Study
We also examine the utility of the contextual features we provide to the prompting module. noPrompt means that we remove all the contextual features from our framework. Others (e.g., Wind Speed) indicate using only one contextual feature. The results on MAE and RMSE are presented in Figure 7. In the absence of any contextual features, we separately add one feature at each time to assess its contribution. Here noPrompt indicates the performance of NRFormer without utilizing any additional contextual features. We can observe that incorporating each feature consistently improves the model performance beyond the noPrompt. In summary, the contextual features have a significant impact on nuclear radiation, and taking them into account is beneficial for forecasting.
@@ 5.3. Ablation Study
5.4. Deployed System Evaluation
To evaluate the real-world deployment performance of the NRFormer framework, we conducted a rigorous temporal generalization analysis using the most recent radiation monitoring data. As shown in Figure 8, the system maintained robust predictive capabilities during the critical evaluation window of January 25–31, 2025, achieving high accuracy in seven-day rolling averages despite significant temporal distributional shifts between the training period (March 17, 2021 – January 1, 2024) and the testing phase. This performance demonstrates the system’s strong stability and robustness, thereby highlighting its practical utility for both individual users and government agencies by empowering them to make timely, data-driven decisions regarding emergency response and public safety.
6. Related Works
Transformer. The advent of the Transformer architecture (Vaswani et al., 2017) has been making a transformative impact across various fields, particularly in time series forecasting (Zhou et al., 2021; Cao et al., 2023; Chowdhury et al., 2022; Gorbett et al., 2023; Wu et al., 2022a; Wang et al., 2023a; Jia et al., 2024) and graph mining (Zhang et al., 2021; Zhu et al., 2023; Fan et al., 2021; Li et al., 2022; Wu et al., 2022b; Chen et al., 2021; Wang et al., 2023b). Transformer depends on self-attention mechanism to capture data correlations and is capable of modeling complicated and long-range dependencies. Due to the powerful capability, recently considerable endeavors have been made to adapt Transformer to time series and spatio-temporal data (Zha et al., 2021; Feng and Tassiulas, 2022; Liang et al., 2023; Zhang and Yan, 2022; Liu et al., 2023; Li et al., 2023). For instance, by introducing a ProbSparse self-attention mechanism, Informer (Zhou et al., 2021) enhances the Transformer framework for time series forecasting with high efficiency and scalability. As another example, Autoformer(Wu et al., 2021) further pushes the boundaries of time series forecasting by integrating decomposition and auto-correlation mechanism into Transformer. In this paper, we adapt the Transformer model with tailored designs to address the unique challenges in nuclear radiation forecasting.
Figure 7. Ablation study of each contextual feature.
@@ 6. Related Works
Figure 8. Performance evaluation of deployed radiation forecasting system using recent data from January 25-31, 2025.
@@ 7. Conclusion
In this paper, we introduce NRFormer, a spatio-temporal graph Transformer for national-wide nuclear radiation forecasting. Specifically, we first devise a non-stationary temporal attention module to capture the erratic and non-stationary characteristics in radiation time series. Subsequently, we propose an imbalance-aware spatial attention module to address the imbalanced spatial distribution issue by adaptively re-weighting the influence of stations from both macro-scale and proximity-constrained perspectives. Furthermore, a radiation propagation prompting module is developed to guide the predictive modeling process. Extensive experiments on real-world nuclear radiation datasets demonstrate the superiority of NRFormer against seven baselines. In the future, we plan to deploy NRFormer to more countries so as to empower various decision-making tasks such as emergency response planning, thereby protecting environmental safety and public health.
Appendix A Supplementary of Experiment
A.1. Baselines
@@
  • DCRNN (Li et al., 2018): a spatio-temporal forecasting model that captures spatial dependencies through bidirectional random walks, while addressing temporal dependencies using encoder-decoder architecture with scheduled sampling.
@@
  • PatchTST (Nie et al., 2023): PatchTST is a Transformer-based model for multivariate time series forecasting, leveraging subseries patching and channel independence to reduce complexity and improve long-term accuracy, also excelling in self-supervised tasks and setting new accuracy benchmarks.
  • Koopa (Liu et al., 2024): Koopa introduces a novel approach for forecasting non-stationary time series by leveraging modern Koopman theory to disentangle and advance time-variant and time-invariant dynamics.
A.2. Evaluation Metrics
We utilize three commonly adopted metrics to evaluate the performance of the model, namely Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are defined as follows. MAE, by contrast, measures the average magnitude of errors without considering their direction, offering a clear and straightforward interpretation of overall accuracy:
(12) $MAE=\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}|\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij}|.$
RMSE calculates the square root of the average squared differences between predicted and observed values, thereby placing greater emphasis on larger errors:
(13) $RMSE=\sqrt{\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}(\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij})^{2}}.$
Meanwhile, MAPE expresses these errors as a percentage of the actual values, facilitating easier comparisons across different scales and datasets. Together, these metrics provide a comprehensive evaluation of the model’s predictive capabilities and robustness:
@@ A.2. Evaluation Metrics
Figure 9. Parameter sensitivity of NRFormer.
[…]
A.3. Parameter Sensitivity Analysis
Figure 10. The deployed web-based radiation monitoring and forecasting system. Left panel: Forecasts for individual radiation monitoring stations, with interactive stations enabling users to access detailed predictions for the coming days. Right panel: Radiation dispersion dynamics across Japan, highlighting spatiotemporal propagation patterns over the next several days.
[…]
Figure 11. The data visualization of nuclear radiation in Japan: the left panel shows the distribution of radiation and meteorological monitoring stations; the top-center chart illustrates the average daily radiation levels across different prefectures; the top-right chart depicts the number of radiation monitoring stations in each prefecture; and the bottom-right chart presents temporal radiation patterns observed across various prefectures.
[…]